OpenStudy (anonymous):

x^2-14x-4=0, solve by completing the square

OpenStudy (amistre64):

hmmm, lets approach this from the perfect (complete) square concept

OpenStudy (amistre64):

(x+a)^2 is a perfect square (x+a)^2 = x^2 + 2a + a^2 given: x^2-14x-4 when does 2a = -14?

OpenStudy (amistre64):

and therefore, what is a^2? to complete the square, we need to add a useful form of zero to this. (a^2-a^2) = 0

OpenStudy (anonymous):

ok i think i have figured this site out now.. these equations are blowing my mind

OpenStudy (amistre64):

they do get easier after some practice :)

OpenStudy (amistre64):

except for the fraction ones ... those are never easy to most people

OpenStudy (anonymous):

lord i hope so..i was getting throught them fine and bam i'm like a deer in the headlgihts

OpenStudy (amistre64):

so, following my setup, what do we get for "a" fromthis? 2a = -14

OpenStudy (anonymous):

a=-7

OpenStudy (amistre64):

good, "a" will always be half the "x" term of the poly given what is a^2 knowing that a = -7?

OpenStudy (anonymous):

ok so where did the 2 come in

OpenStudy (anonymous):

is it from the 4???

OpenStudy (amistre64):

a "complete square" is what we are trying to work out here. A complete square is a perfect square. perfect squares are: 1^2, 2^2, 3^2 ... if using numbers if using a binomial of the form (x+a) then we just square it: (x+a)^2 is a perfect square

OpenStudy (amistre64):

(x+a)^2 = (x+a) (x+a) = x^2 +2a +a^2 ; if you wanna use "foil"

OpenStudy (amistre64):

this gives a general setup to compare with the specific poly given: x^2 -14x +a^2 -a^2 - 4 x^2 +2a +a^2 + _______ in order to "complete the square" we add and subtract a^2 to the given poly

OpenStudy (amistre64):

since a is -7, lets add and subtract (-7)^2 = 49 from from it; why? (becasue we dont want to change the value of the poly, just the way it looks; so adding zero helps us out) x^2 -14x +a^2 -a^2 - 4 x^2 -14x +49 - 49 - 4 group the perfect square (x^2 -14x +49) - 49 - 4 and rewrite it in its (x+a)^2 form (x-7)^2 - 53 = 0

OpenStudy (anonymous):

ok so since x-7 in brackets it stays x-7 and i add 53 to both sides right

OpenStudy (amistre64):

you can yes

OpenStudy (anonymous):

ok whats the easiest and simple way to do this

OpenStudy (amistre64):

well, completing the square is the proof to the quadratic formula

OpenStudy (anonymous):

what is i make 53 a sq root and it would became 7sqrt53

OpenStudy (anonymous):

*if

OpenStudy (anonymous):

7 +/-sqrt53 right?

OpenStudy (anonymous):

ugh i've confuesd myself

OpenStudy (amistre64):

\[ax^2+bx+c = 0\] \[x^2+\frac bax+\frac ca = 0\] \[x^2+\frac bax+(\left(\frac b{2a}\right)^2-\left(\frac b{2a}\right)^2)+\frac ca = 0\] \[\left(x^2+\frac bax+(\frac b{2a})^2\right)^2-\left(\frac b{2a}\right)^2+\frac ca = 0\] \[\left(x+\frac b{2a}\right)^2-\left(\frac b{2a}\right)^2+\frac ca = 0\] \[\left(x+\frac b{2a}\right)^2-\frac {b^2}{4a^2}+\frac ca = 0\] \[\left(x+\frac b{2a}\right)^2-\frac {b^2}{4a^2}+\frac {4ac}{4a^2} = 0\] \[\left(x+\frac b{2a}\right)^2-\frac {b^2}{4a^2}+\frac {4ac}{4a^2} = 0\] \[\left(x+\frac b{2a}\right)^2=\frac {b^2}{4a^2}-\frac {4ac}{4a^2} = 0\] \[\left(x+\frac b{2a}\right)^2=\frac {b^2-4ac}{4a^2}\] \[x+\frac b{2a}=\pm\sqrt{\frac {b^2-4ac}{4a^2}}\] \[x=-\frac b{2a}\pm\sqrt{\frac {b^2-4ac}{4a^2}}\] \[x=-\frac b{2a}\pm\frac {\sqrt{b^2-4ac}}{2a}\] \[x=-\frac {-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (amistre64):

that is correct: x = 7 +/-sqrt53 right

OpenStudy (amistre64):

dont confuse the "a" in the formula with the random letter i choose to begin the post with :)

OpenStudy (anonymous):

it kicked it out on my internet college math cours, it threw up up the \[\left(\begin{matrix}b \\ 2\end{matrix}\right)^2\]

OpenStudy (amistre64):

ax^2 + bx + c = 0 x^2 -14x -4 = 0 \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[x=\frac{-(-14)\pm\sqrt{(-14)^2-4(1)(-4)}}{2(1)}\] \[x=\frac{14\pm\sqrt{196+16}}{2}\] \[x=\frac{14\pm\sqrt{212}}{2}\] \[x=\frac{14\pm\sqrt{4*53}}{2}\] \[x=\frac{14\pm2\sqrt{53}}{2}\] \[x=7\pm\sqrt{53}\]

OpenStudy (anonymous):

ah it is wanting two numbers instead of a sqrt

OpenStudy (amistre64):

hmm, yeah, i have never been to good with formating things the way a program wants it formated :)

OpenStudy (anonymous):

the example shows x=11+/- 5 sqrt5

OpenStudy (amistre64):

if you can take a screenshot of the problem, i could get a glimpse of what it is expecting an answer to look like :)

OpenStudy (anonymous):

i tried to copy and paste it and it won't let me i just sent an email to my professior, i'm waiting on her reply..thank you for your help tho. i'll send message once i talk to her

OpenStudy (amistre64):

ok, good luck ;)